OpenStudy (anonymous):

What are the zeros of 2x^2-4x-14

OpenStudy (anonymous):

You can calculate the zeros by taking the factors of the first term and the last term (1,2,7,14) and use synthetic division to figure it out. Or factor it :p There is really a lot of ways to go about it.

OpenStudy (anonymous):

I factored it to 2(x^2-2x-7) but idk where to go from there :(

OpenStudy (anonymous):

You continue factoring?

OpenStudy (zzr0ck3r):

\(2x^2-4x-14=0\implies x^2-2x-7=0\\(x-1)^2-7-1=0\\(x-1)^2=8\\x-1=\pm \sqrt{8}=\pm2\sqrt{2}\\x=\pm 2 \sqrt{2}+1\)

OpenStudy (anonymous):

Well, there you have it ^. But learn to do it...it's easy. Or use wolfram alpha.

OpenStudy (anonymous):

Thank you!!

OpenStudy (zzr0ck3r):

for these ones when you cant use the a,c method, you can use the quadratic formula or you can complete the square. I suggest learning to complete the square because it will always work and is not that hard.

OpenStudy (zzr0ck3r):

I just took half the middle term "-1" and then you do (x+(-1))^2-7-(-1)^2 = (x-1)^2-8